• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 19:11:28 on October 24, 2010

    Entry number 336539

    keyword=Run plan for Sunday swing to Monday Day

    We have been pretty steadily taking production data at 4 pass for the
    last couple days. Friday we changed kinematics to the low energy, Q2=2
    point and have been accumulating statistics steadily. 
    We are still in a mode in which we are triggering on the HRS only, but
    steady progress is being made on the calo trigger.
    Tonight, we will continue to take production data (see notes on white
    board for up-to-date run conditions- current, CODA config. etc.)
    Tomorrow (Monday) morning, we will change HRS polarity to positive in
    order to take more elastic data. Since the rates relatively high, we may
    get 2 iterations of calorimeter gain matching before Tuesday.
    Tuesday we will change HRS polarity back to negative during the planned
    beam studies and return to production data taking.
    Longer term plans:
    1. Moller on Wednesday DAY shift.
    2. We may want to do an Arc energy measurement at 4 pass. TBD.
    Things to keep an eye on:
    1. The pockels cell tripped off for a while last week (i.e. beam was not
    polarized). We should keep an eye on this.
    2. The beam current seems to drift a little - we need to maintain at 2 uA
    - dead time will increase rapidly if the current creeps up.
    3. Hall C is starting to get their beam modulation (including beam energy
    modulation) system going. We need to make sure we have the appropriate
    signals in our DAQ that tell us when the energy lock is off, and when the
    modulation is ongoing.
    4. Compton rates are still high.

    A copy of this log entry has been emailed to: jroche@jlab.org,camsonne@jlab.org,folts@jlab.org,hyde@jlab.org,munoz@jlab.org,glamazdi@jlab.org